OpenStudy (technodynamic):

A nitric solution containing 308g of HNO3 per liter of solution has a density of of 1.35g/cm3. 1. Calculate the mass percent of HNO3 in this solution. 2. Calculate the molality of HNO3 in this solution. I already know that the mass%=mass solute/mass solution. I'm confused how to i set up the density part? Please explain step by step. I have to study this for an upcoming test. I will reward you with a trophy.

OpenStudy (dtan5457):

Look, I'm pretty crappy with chemistry but I'm trying to learn it, I did a ton of these questions yesterday. 1.I'm pretty sure what this means is that 308g is your solute, but your solution will be determined from the density. d=m/v v=1000g(1 liter) d=1.35g/cm mass=d x v mass=1350 grams or 1.35 liters your formula of mass%=308/1350 x 100 I got 22.8%. I'm pretty sure this is right

OpenStudy (dtan5457):

As for the 2nd one, molarity is #of moles/liters that means you need to first calculate the amount of moles, which is given weight/formula weight HNO3 given weight=308gram h=1 n=14 o3=48 formula weight=63 308/63=roughly 4.8 moles 4.8/solution solution is 1 liter since i'm pretty sure the molarity formula HAS to be in liters 4.8 moles, 1 liter 4.8/1=4.8 molarity

OpenStudy (dtan5457):

i hope i'm right lol..

OpenStudy (dtan5457):

@JFraser can you check this for me?

OpenStudy (technodynamic):

Yes you're correct! Thank you!!

OpenStudy (dtan5457):

LOL I am??? This might be the first chemistry question I did correctly in 6 months, that's cool..

OpenStudy (dtan5457):

If you have anymore of these, I'll gladly look at them as I need the practice too..

OpenStudy (technodynamic):

Lol yeah chem can be crazy hard. The exam is this Friday too, I'm glad you showed the steps because I wasn't sure if I was solving it right. Thanks for clearing up my confusion. Lol

OpenStudy (dtan5457):

Yeah, gas laws are especially a pain to remember. (charles,boyles etc) It may take a lot of practice, also requires critical reading skills...